Part A The QuickFixCompany just paid a dividend of $1.25 and analysts expect the dividend to...

Part A

The QuickFixCompany just paid a dividend of $1.25 and analysts expect the dividend to grow at its compound average growth rate of 10.72% forever.

If you plan on holding the stock for just 7 years, and you have an expected rate of return of 14%, how much would you pay for the stock?

Assume that the next owner also expects to earn 14% on his or her investment.

Show your work in the form of the calculator functions you use and the numbers you input into the calculator, and provide a brief explanation of what you have done/steps you have taken and why.

Part B

As a follow up to Assignment #1 (see the course home page): Assume it is 7 years later and the Investor who purchased the stock 7 years ago, price of which you calculated in Assignment 1, is ready to sell the stock now.

What would be the price he can sell it for, assuming the grown rate of dividends has remained at 10.72% forever and the required rate of return has remained at 14%?

Show your work and provide a brief explanation of what you have done/steps you have taken and why.

I ONLY NEED HELP WITH PART B!
Answer #1

Dividend in year 8 =D0*(1+growth)^8 =1.25*(1+10.72%)^8 =2.8230504
Required rate =14%
Price of the stock using dividend discount model =D8/(Required-Growth) =2.8230504/(14%-10.72%) =86.07